You may need to draw the figures indicated on your scratch paper to answer this question.
Write the equation in standard form of the line that passes through P (5, - 7) and has slope of -1 over 5.
The equation is:

Answer by stanbon(57407)

Write the equation in standard form of the line that passes through P(5,-7) and has slope of -1 over 5.
---
Form:
y = mx + b
Substitute for y,m,and x; solve for "b":
-7 = (-1/5)5 + b
-7 = -1 + b
b = -6

The equation is:
y = (-1/5)x-6
-----------------
Cheers,
Stan H.